# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v1

@@ 5.2 Real-world Results
In the real-world tasks, we employ diffusion policy (DP) chi2023diffusionpolicy and ACT zhao2023act to drive robots in performing real-world tasks, where the visual encoders are pre-trained using MAE and R3M. Policy training details are presented in Appendix  C.
To evaluate the effectiveness and generalization of H2R, we conduct real-world experiments under two different robot embodiments: UR5 (same embodiment as downstream execution) / Franka (cross-embodiment generalization).
To assess the effectiveness of H2R, we use H2R-UR5-SSv2-1M-Mix during pre-training, which matches the embodiment used in policy training and evaluation. As shown in Table 3, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies. For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M), while for Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M). Similar trends are observed under ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), and Gripper tasks improving by 8.3% and 11.7%, respectively.
To evaluate cross-embodiment generalization, we use H2R-Franka-SSv2-1M dataset. Downstream policy training and evaluation are still conducted with UR5 Leaphand. As shown in Table 6, H2R still outperforms raw Ego4D despite the embodiment mismatch. Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%. Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%. Although Franka-based pretraining underperforms UR5, the results confirm H2R’s robustness to embodiment variations. Additional analysis, including failure case categorization and experiments on generalization under lighting perturbations, is provided in Appendix  E and  G.
@@ 5.3 Performance on Other Datasets
We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented H2R-UR5-Ego4D-1M-Mix, and apply them to real-world Leaphand manipulation tasks using the ACT zhao2023act policy. As shown in Table 5, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M. These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
Table 7: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget). Task success rates (%) are reported under DP and ACT policies. Removing either component causes significant performance degradation, with red values showing the relative drop compared to full H2R.
@@ 6 Conclusion
We propose H2R, a data augmentation technique that bridges the visual gap between human hand demonstrations and robotic arm manipulations by replacing human hands in first-person videos with robotic arm movements. Using 3D hand reconstruction and image inpainting models, H2R generates synthetic robotic arm manipulation sequences, making them more suitable for robot pre-training. Experiments across simulation benchmarks and real-world tasks demonstrate consistent improvements in success rates for encoders trained with various pre-training methods (e.g., MAE, R3M), highlighting the effectiveness and generalization of H2R. H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection. Finally, we expressed the broader impact of this work in Appendix H.
Appendix A Limitations
H2R is a simple data augmentation technique for robot pre-training from videos. There are several limitations, which open up possibilities for future work:
  1. 1.
    H2R augmentation pipeline is sensitive to the accuracy of pose estimation and hand segmentaion, which demands high-quality egocentric video datasets as prerequisites. This sensitivity becomes particularly pronounced in challenging scenarios involving low-light conditions or partial hand occlusions, where current estimation methods often fail to maintain required precision.
  2. 2.
    While our robotic platform incorporates diverse manipulator configurations spanning various robotic arm, grippers and dexterous hands, the current implementation remains limited to single-arm operation. Future work will expand this framework to dual-arm and humanoid robotics, significantly broadening the model’s application scope.
  3. 3.
    It would be valuable to extend our work on more tasks besides manipulation, like visual navigation and mobile manipulation.
Appendix B Details of Simulator Camera Position Alignment
@@
Appendix C Policy Training Details
Policy Training in Simulation Experiment. For Robomimic tasks, we train for 200 steps and report the mean success rate. For tasks in RLBench, we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization. For the PushT task, we train the Diffusion Policy model for 200 epoches and report the success rate in the simulation environment. The training hyperparameters used in this work are identical to those described in the original paper.
Policy Training in Real-world Experiment. We select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section  4.1. DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
Appendix D Evaluation of H2R Effectiveness
In Figure 7 are six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 7 (left to right) are described as follows:
  • holding a yellow measuring tape at a construction site
  • holding a bottle
  • painting on the wall with a brush / Brushing paint on a wall
  • turning on the kitchen faucet
  • preparing painting materials on a table
The similarity scores computed by CLIP are presented in Table 8. Across these samples, the augmented images consistently exhibit higher similarity scores with the robot-centric action descriptions compared to the similarity scores between the unaugmented images and the human-centric action descriptions. This quantitative analysis demonstrates that the H2R augmentation effectively enhances the visual alignment between human and robotic actions.
Img1 Img2 Img3 Img4 Img5 Img6
ori 30.6 23.7 31.2 31.7 27.5 28.7
aug 32.6 28.9 32.3 32.0 28.7 31.0
@@ Appendix D Evaluation of H2R Effectiveness
Appendix E Failure Case Analysis
@@
Appendix F Additional Ablation Study in Simulation Experiment
@@
Appendix G Generalization on Light Condition.
@@
Appendix H Broader Impact
In this work, we proposed a a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric video. We focused on the theme above and doesn’t have direct social impact. We hope that through our data augmentation methods, more egocentric human datasets can be used for higher-level robot manipulation models, thereby promoting progress in the field of embodied intelligence.

### v3

@@ 5.1 Real-world Results
In the real-world tasks, we employ diffusion policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] to drive robots in performing manipulation tasks, where the visual encoders are pre-trained using MAE and R3M. To evaluate the effectiveness and generalization of H2R, we first conduct experiments to validate its performance enhancement when the pre-training and policy training embodiments are the same. We then perform cross-embodiment experiments to investigate the transfer effects of H2R when the pre-training and policy training embodiments differ. We finally apply H2R to fine-tune a pre-trained vision backbone in a Vision-Language-Action model, showcasing its promising potential in this advanced setting.
For identical embodiment experiment, we use UR5 or Franka robot for augmentation during pre-training, which matches the embodiment used in policy training and evaluation. As shown in Table 6, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies. For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M). For Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M). For Dual-arm Franka tasks, the improvements reach 3.3% (MAE) and 7.3% (R3M). Similar trends are observed in ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), Gripper tasks improving by 8.3%(MAE) and 11.7%(R3M), and Dual-arm Franka tasks improving by 16.7%(MAE) and 5%(R3M).
To evaluate cross-embodiment generalization, we use Franka-augmented the SSv2 dataset with H2R pipeline. Downstream policy training and evaluation are conducted with UR5 Leaphand. As shown in Table 7, H2R still outperforms raw Ego4D despite the embodiment mismatch. Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%. Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%.
@@ 5.2 Performance on Other Datasets
We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented dataset, and apply them to real-world Leaphand manipulation tasks using the ACT [zhao2023act] policy. As shown in Table 10, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M. These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
5.3 Potential for VLA Model Pretraining
Figure 5: Dual-UR5 tasks. The left columns shows PlaceBowl task and the right columns shows StackBlocks task.
A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs). In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning. Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [li2025unifiedvideoactionmodel] separately on both the original SSv2 and the H2R-augmented datasets. The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5 robot tasks(shown in Figure  5). The results shown in Table  11 demonstrate a clear performance trend in the different variants of the UVA model.
Table 11: Success rates of UVA variants with different visual backbones.
Model PlaceBowl StackBlocks Average UVA 0.2 0.2 0.2 UVA + SSv2 0.1 0.2 0.15 UVA + SSv2 + H2R 0.4 0.3 0.35
@@ 6 Conclusion
We propose H2R, a data augmentation technique that bridges the visual gap between human hand demonstrations and robotic arm manipulations by replacing human hands in first-person videos with robotic arm movements. Using 3D hand reconstruction and image inpainting models, H2R generates synthetic robotic arm manipulation sequences, making them more suitable for robot pre-training. Experiments across simulation benchmarks and real-world tasks demonstrate consistent improvements in success rates for encoders trained with various pre-training methods (e.g., MAE, R3M), highlighting the effectiveness and generalization of H2R. H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection.
[…]
Supplementary Material
Appendix A Details of Simulator Camera Position Alignment
@@
Appendix B Details of Experimental Setup
Figure 6: Simulation benchmark. We choose 3 tasks from the Robomimic, 3 tasks from the RLBench, and 3 tasks from the CortexBench, covering a range of robotic manipulation skills. We also include the PushT task, designed for the Diffusion Policy framework, as an additional benchmark to evaluate performance in a different task setup.
Figure 7: Real-world Robot Setup. Illustration of different real-world experimental setup.
Data Collection in Real-world Experiment. For the UR5-gripper, demonstration trajectories were collected via keyboard-based teleoperation. For the UR5-Leaphand, demonstrations were gathered using vision-guided teleoperation. For the Dual-arm Franka and Dual-arm UR5 setups, data collection was performed through homogeneous-arm bilateral teleoperation.
Additional Evaluation Details. Figure  6 provides visualizations of the different simulation benchmarks, and Figure  7 illustrates the real-world experimental setup. For the simulation benchmarks, we use the default configurations for both training and evaluation. In the real-world tasks, the UR5–Gripper and UR5–Leaphand setups use the camera shown in Figure  7 as the visual input, whereas the dual-arm Franka and dual-arm UR5 configurations use a front–top camera together with two wrist-mounted cameras (three cameras in total) for visual observation. To ensure that the initial conditions are evaluated consistently across algorithms, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstration collection. During evaluation, the object placement region is kept identical to that used in data collection, ensuring a fair and consistent comparison.
@@ Appendix B Details of Experimental Setup
Appendix C Failure Case Analysis
@@
Appendix D Additional Ablation Study in Simulation Experiment
@@
Appendix E Generalization on Light Condition.